OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

First you need to distribute. Then combine like terms.

OpenStudy (anonymous):

i did and i got the wrong answer

OpenStudy (anonymous):

3x+3=-2x+1+6 Did you distribute correctly?

OpenStudy (anonymous):

no i messed up. can you help with another few questions?

OpenStudy (anonymous):

One or two...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4.

OpenStudy (anonymous):

Let's see you distribute first

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

-2x+1=-2x-1-4

OpenStudy (anonymous):

thats what i got so far

OpenStudy (ranga):

In the fifth reply the right hand side should be: -2x + 2 + 6

OpenStudy (anonymous):

Okay, so your problem is that you're not distributing correctly. You need to multiple the number by both terms when distributing

OpenStudy (anonymous):

Ugh, thank you @ranga I hate typos

OpenStudy (anonymous):

i thought i did that

OpenStudy (anonymous):

-2x-6=-2x-2-4

OpenStudy (haseeb96):

3x+3=-2x+2+6 3x+2x=2+6-3 5x=5 x=5\5 x=1 if you want to check answer then put 1 in both sides let i put in it 3(1+1)=-2(x-1)+6 3(2)=-2(1-1)+6 6=2(0)+6 6=6 Hence proved

OpenStudy (anonymous):

ok thamks @Haseeb96